A 3kg toy car with a speed of 6 m/s collides head on with a 2kg car traveling in the opposite direction with a speed of 4m/s. If the cars are locked together after the collision with a speed of 2m/s, how much kinetic energy is lost?
Answer #1

Conserving momentum,

m1*v1 + m2*v2 = (m1+m2)*v

So, 3*6 + 2*(-4) = (3+2)*v

So, v = 2 m/s

Initial KE of the system = 0.5*3*6^2 + 0.5*2*4^2

= 70 J

Final KE of the system:

= 0.5*(3+2)*2^2

= 10J

So, KE lost = 70 - 10 = 60J